Circular No. 8141

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003ew, 2003ex, 2003ey, AND 2003ez Further to IAUC 8140, several additional apparent supernovae have

been reported by L. G. Strolger et al. (2003ew), by W. M. Wood-Vasey et al. (2003ex, 2003ey) and by M. Papenkova and W. Li (2003ez).

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003ew Jan. 4.7 12^{h}36^{m}27.^{s}78 +62^{o}11′25′′.1 24.1 0′′.21 W, 0′′.1 S
2003ex Mar. 26.44 13 27 32.31 + 5 31 14.4 19.1 2′′.5 W, 1′′.1 S
2003ey May 24.33 14 20 03.28 − 8 28 40.0 19.2 2′′.4 W, 0′′.8 S
2003ez June 3.2 12 43 11.87 − 2 30 04.1 17.5 4′′.5 W, 0′′.7 S

Further magnitudes provided by the respective observers: SN 2003ew, 2003 Feb. 20.4 UT, 24.3; May 21.8, [26.0. SN 2003ex, 2002 May 6.23, [21.4; 2003 Feb. 5.57, [20.8; Apr. 6.39, 19.8; 14.40, [19.1; 20.41, 19.8; 24.42, [20.3; 25.59, [19.2; 26.37, 19.8; 28.40, [18.4; 29.39, 20.2; May 2.39, [19.5; 3.34, 19.8; 6.40, [19.8; 7.35, 20.1; 8.29, 20.4 (uncertainty up to 0.4 mag on the NEAT data). SN 2003ey, 2002 Apr. 8.48, [21.4; 2003 Mar. 27.45, [21.1; Apr. 9.57, [19.4; 27.46, [19.5; May 2.42, [19.0; 30.28, 18.8. SN 2003ez in PGC 42782, May 21.2, [19.0; June 4.2, 17.4.

NOVA IN NGC 6822 Papenkova and Li also report the LOTOSS discovery of an apparent

nova in NGC 6822 on unfiltered KAIT images taken on May 28.4 (mag ∼ 19.0) and June 3.4 UT (mag ∼ 19.2). The new object is located at α = 19^{h}45^{m}04^{s}.18, δ = −14^{o}45′46′′.5 (equinox 2000.0), which is 0′′.8 west and 2′′.7 north of a foreground star of mag ∼ 18.5. A KAIT image taken on 2002 Sept. 4.3 showed nothing at this position (limiting mag ∼ 20.0).

SUPERNOVA 2003ep IN NGC 7053 T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center

for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2003ep (cf. IAUC 8140), obtained by P. Berlind on June 4.46 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a reddened type-Ia supernova, 3–4 weeks past maximum. There is a relatively strong (equivalent width ∼ 0.15 nm) Na I D absorption line at the velocity of the host galaxy.

2003 June 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green